# ALIFE2024 template

### v1

@@ Introduction
Simulations afford the opportunity to model individual behaviours and interactions within a controlled setting, enabling the study of emergent properties in a systematic and repeatable manner. Recently, efforts to merge deep learning and complex systems has greatly improved CI research capabilities by enabling simulations to scale to hundreds or thousands of individuals; many more than was previously possible Ha and Tang, (2022). By leveraging real-world data from living systems, we can simulate environments with the same conditions in which collective behaviours naturally occur, allowing for the exploration of factors such as communication, information aggregation (Kameda et al.,, 2022), coarse-graining (Flack,, 2017), and their impacts on the emergence of collective behaviours.
@@
We chose to reproduce the dynamics of the ant species Pristomyrmex punctatus which is characterised by its lack of a queen caste, with mature colonies often consisting of thousands of female workers. Instead of constructing nests, the colony exhibits a strong clustering behaviour (Tsuji and Dobata,, 2011). (For further information on the experimental setup and the detection/tracking pipeline, see Maruyama et al., (2023).) Based on data extracted from video recordings, we built a simulation environment in which the agent (ant) begins each trial at position $P_{\alpha}(0)=P_{\tau}(0)$, the beginning of a (real) target ant trail. Figure 1 shows the comparison between the video recordings and the simulation environment. Currently, we select a target ant at random from the data whilst ensuring that the target moved a modest amount during the $t_{\text{lim}}=30$ second time period. The target ant and corresponding trail are depicted in red, while the agent and agent trail are shown in blue. This is seen more clearly in Figure 2 where we also depict the ant’s vision which we further separate into four segments: $5$ in the forward-facing direction (indicated by the blue line depicting the agent’s direction $\theta$), and $3$ facing backwards. Each ant in the environment begins at some position translated to pixel coordinates $(x,y)$ from the real data, relative to the arena’s dimensions and scale.

### v2

@@ Introduction
Simulations afford the opportunity to model individual behaviours and interactions within a controlled setting, enabling the study of emergent properties in a systematic and repeatable manner. Recently, efforts to merge deep learning and complex systems has greatly improved CI research capabilities by enabling simulations to scale to hundreds or thousands of individuals; many more than was previously possible (Ha and Tang,, 2022). By leveraging real-world data from living systems, we can simulate environments with the same conditions in which collective behaviours naturally occur, allowing for the exploration of factors such as communication, information aggregation (Kameda et al.,, 2022), coarse-graining (Flack,, 2017), and their impacts on the emergence of collective behaviours.
@@
We chose to reproduce the dynamics of the ant species Pristomyrmex punctatus which is characterised by its lack of a queen caste, with mature colonies often consisting of thousands of female workers. Instead of constructing nests, the colony exhibits a strong clustering behaviour (Tsuji and Dobata,, 2011). (For further information on the experimental setup and the detection/tracking pipeline, see Maruyama et al., (2023).) Based on data extracted from video recordings, we built a simulation environment in which the agent (ant) begins each trial at position $P_{\alpha}(0)=P_{\tau}(0)$, the beginning of a (real) target ant trail. Figure 1 shows the comparison between the video recordings and the simulation environment. Currently, we select a target ant at random from the data whilst ensuring that the target moved a modest amount during the $t_{\text{lim}}=30$ second time period. The target ant and corresponding trail are depicted in red, while the agent and agent trail are shown in blue. This is seen more clearly in Figure 2 where we also depict the ant’s vision which we further separate into eight segments: $5$ in the forward-facing direction (indicated by the blue line depicting the agent’s direction $\theta$), and $3$ facing backwards. Each ant in the environment begins at some position translated to pixel coordinates $(x,y)$ from the real data, relative to the arena’s dimensions and scale.